A projectile that is fired vertically into the air with an initial velocity of 120 ft per second can be modeled by the equation s=120t-16t2. In the equation, s is the distance in feet of the projectile above the ground after t seconds. Is it possible for the projectie to reach 900 feet? Justify you answer.
Answer by stanbon(75887)

A projectile that is fired vertically into the air with an initial velocity of 120 ft per second can be modeled by the equation s=120t-16t2. In the equation, s is the distance in feet of the projectile above the ground after t seconds. Is it possible for the projectie to reach 900 feet? Justify you answer.
------------------
See if there is a Real Number solution for:::
-16t^2 + 120t = 900
---
b^2-4ac = 120^2-4*-16*-900 is less than zero
-----
So t is a complex number.
So the projectile never reaches 900 ft.
-----
Cheers,
Stan H.